# StoryTeller: Improving Long Video Description through Global Audio-Visual Character Identification

### v2

@@ Abstract
[…]
Existing large vision-language models (LVLMs) are largely limited to processing short, seconds-long videos and struggle with generating coherent descriptions for extended video spanning minutes or more. Long video description introduces new challenges, such as consistent character identification and plot-level descriptions incorporating both visual and audio information. To address these, we figure out audio-visual character identification, matching character names to each dialogue, as a key factor. We propose StoryTeller, a system for generating dense descriptions of long videos, incorporating both low-level visual concepts and high-level plot information. StoryTeller uses a multimodal large language model that integrates visual, audio, and text modalities to perform audio-visual character identification on minute-long video clips. The results are then fed into a LVLM to enhance consistency of video description. We validate our approach on movie description tasks and introduce MovieStory101, a dataset with dense descriptions for three-minute movie clips. To evaluate long video descriptions, we create StoryQA, a large set of multiple-choice questions for MovieStory101 test set. We assess descriptions by inputting them into GPT-4 to answer these questions, using accuracy as an automatic evaluation metric. Experiments show that StoryTeller outperforms all open and closed-source baselines on StoryQA, achieving 9.5% higher accuracy than the strongest baseline, Gemini-1.5-pro, and demonstrating a +15.56% advantage in human side-by-side evaluations. Additionally, incorporating audio-visual character identification from StoryTeller improves the performance of all video description models, with Gemini-1.5-pro and GPT-4o showing relative improvement of 5.5% and 13.0%, respectively, in accuracy on StoryQA. Datasets and code are available at https://github.com/hyc2026/StoryTeller.
@@ 1 Introduction
Generating detailed video descriptions is a fundamental challenge in video understanding. While large vision-language models (LVLMs) have made significant progress [45, 27, 8, 44, 26, 25], they remain limited to processing seconds-long videos and generating low-level visual concepts like objects, scenes and atomic actions. Real-world videos, such as movies, are much longer and require high-level information for effective description, such as plot development. Recent research works address long videos by segmenting them, generating descriptions for each segment, and combining them using large language models [51, 13, 14]. However, there is still a gap to resolve the key challenges of long video description: maintaining consistent character identification and incorporating dialogue, which conveys rich plot information.
Thus, accurate character identification and tracking across both visual and audio modalities are essential for high-quality long video descriptions. While existing studies [35, 20] have improved the quality of video descriptions by focusing on character identification and tracking through visual cues, a critical gap remains in linking dialogue with on-screen characters to identify speakers. Bridging this gap is essential for enhancing long video descriptions.
Figure 1: Pipeline of StoryTeller, consisting of three main modules: Stage I - Video Segmentation, where long videos are divided into seconds-long clips that are relatively independent and internally complete; Stage II - Audio-Visual Character Identification, where characters are identified for each dialogue line using both audio and visual cues; and Stage III - Description Generation, where detailed descriptions are generated for each short clip, ultimately producing a coherent and consistent narrative for the entire long video.
@@
A major challenge in long video description is the lack of training and test data. To address this, we introduce MovieStory101, a new dataset focused on movies—long videos with high information density. The videos in MovieStory101 are from Movie101[49] and Movie101v2 [50], comprising 187 movies split into 5,982 three-minute clips. Each clip is annotated with storyline descriptions, subtitles, and character identity labels for each dialogue line. The storyline annotations ensure a complete understanding of the movie’s plot. By training on this dataset, we aim to unleash the model’s ability to generate plot-level descriptions. In addition, evaluating long video descriptions is also challenging. Besides human evaluation, we propose an automated evaluation method for MovieStory101’s test set: StoryQA. StoryQA has 11 multiple-choice questions in average for each video, covering visual actions, character relationships, and plot details. We evaluate a long video description by inputting it into GPT-4 [31] to answer these questions, using the accuracy as an evaluation metric.
Based on StoryQA and human side-by-side evaluation, we demonstrate that StoryTeller outperforms Gemini-1.5-pro [38], GPT-4o [32], and several advanced open-source LVLMs [27, 23, 45, 7] in generating descriptions for long videos. Specifically, the 7B model-based StoryTeller achieves an accuracy in the StoryQA evaluation that is 9.5% higher than the best-performing baseline model, Gemini-1.5-pro. In human side-by-side evaluation, StoryTeller shows an advantage of +15.56% over Gemini-1.5-pro. Additionally, we verify the significant role of accurate audio-visual character identification on improving description. Incorporating audio-visual character identification results from StoryTeller as input consistently improves the performance of various LVLMs, with Gemini-1.5-pro and GPT-4o achieving relative accuracy gains of 5.5% and 13.0% on StoryQA, respectively.
@@ 2 Related Work
Long Video Description. Limited by the number of input frames, LVLMs are primarily effective for short video clips and typically generate descriptions focused on low-level visual concepts [45, 27, 8, 44, 26, 25]. There remains a gap in describing longer videos, which may span several minutes or more, and in capturing higher-level content like character interactions and plot progression.Some research [51, 13, 14] treats long videos as a series of short clips, incorporating previous descriptions when describing the current one, but this can lead to cumulative errors that affect consistency and coherence of the overall narrative [13]. Another approach, Video ReCap [19], offers a recursive model for varying video lengths. Nevertheless, for longer videos, this method tends to produce summaries, limiting its ability to generate detailed descriptions.
@@
Audio-Visual Character Identification. Character identification is essential for high-quality video descriptions [35, 20], particularly to maintain consistency across clips. While prior research has focused on visual-level identification [30, 41], leveraging audio cues to determine which character is speaking can straightforwardly improve descriptions. A closely related area of research is audio-visual diarization, which integrates both visual and auditory features to segment speakers in audio [46, 16]. Nonetheless, current methods are still unable to identify the speakers’ identities.
3 StoryTeller Long Video Description System
3.1 Overview
3.2 Audio-Visual Character Identification
The audio-visual character identification module determines speakers identity in video clips by integrating audio, visual and textual information. In particular, given a video clip (containing frames, audio and subtitles) and a cast list (with character photos and names), the module assigns each dialogue line to its corresponding character. If the speaker is in the cast list, their name is provided; otherwise, a descriptive identity (e.g., ‘a police officer’ or ‘a little girl’) is used. This task presents several challenges: 1) relying solely on visual information is often insufficient, especially when multiple characters are on screen or the speaker is not visible; 2) audio cues in seconds-long clips are typically isolated, making them less effective for character identification. However, in longer segments, audio can be leveraged to extract speaker features across different clips, such as identifying if the same person speaks in multiple segments. Models that are limited to short segments often struggle to incorporate such global features effectively.
To address the aforementioned challenges, we design a novel audio-visual character identification mechanism incorporating a MLLM equipped with both audio and video modalities. We also develop a global decoding algorithm to leverage global audio information for more accurate character identification. Specifically, as illustrated in Figure 1, we begin by performing audio diarization on the entire video, segmenting the dialogue, and assigning IDs to different speakers (e.g., C1, C2, etc.), as detailed in Appendix D. The subtitles with global IDs are then fed into the MLLM to identify the corresponding character names for each global ID. The global decoding mechanism processes short video clips in parallel, ensuring that global IDs in different clips are consistently mapped to the same name, thereby improving identification accuracy. Details on the implementation of the MLLM and the global decoding algorithm are provided below.
Model Architecture and Training. We integrate a MLLM capable of processing both visual and auditory data, as depicted in Figure 2. For visual inputs, each image or video frame is encoded by CLIP-ViT [36], followed by mapping into the token embedding space of the downstream LLM through a multi-layer perceptron (MLP). These tokens are then fed into the LLM. The audio inputs are processed by dual auditory encoders [40], a speech encoder from Whisper-Large-v2 [37] and a non-speech BEATs audio encoder [4]. The outputs from the two audio encoders are concatenated at the frame level. Subsequently, a window-level Q-former [24] align the combined audio representation to the LLM’s input space.
We initialize the visual module and language model from the Tarsier-7b [44], a large vision-language model demonstrating SOTA performance on various video understanding benchmarks. Our training protocol consists of three phases: (1) pre-training the audio module, (2) fine-tuning on tasks including audio diarization, character identification, recognition and tracking, and (3) fine-tuning for audio-visual character identification using MovieStory101 training set. Details on the dataset used in each phase and the complete training methodology are provided in Appendix E.
Global Decoding. We use the MLLM to assign the names to each global character ID. First, these names come from a set $\mathcal{N}$, which includes all names from the cast list as well as a special item “Others”, representing characters not in the cast list. When the model outputs “Others”, it will further provide a descriptive name. Second, the model outputs a JSON dictionary, where each key is a global character ID within the input short clip, and each corresponding value is the assigned name. Keys are listed in ascending order of global character IDs.
$p_{\theta}(y|v_{k},x)=\prod_{t}p_{\theta}(y_{t}|v_{k},x,y_{<t}),$ (1)
Then, the global probability of $x$ corresponding to name $y$ is defined as
$p^{\rm global}_{\theta}(y|x)=\prod_{k=1}^{m}p_{\theta}(y|v_{k},x).$ (2)
For example, as shown in Figure 2, both videos $v_{1}$ and $v_{2}$ contain lines from character $C1$. In $v_{2}$, although Xu Tailang is the one speaking, he is shown in profile while the camera focuses entirely on Director Jin, who is not speaking. As a result, the model assigns a higher probability to Director Jin being the speaker. However, the model still has over 30% confidence in identifying the speaker as Xu Tailang. By integrating information from the entire video globally, $C1$ is ultimately mapped to Xu Tailang, correcting the erroneous prediction in $v_{2}$.
3.3 Description Generation
After identifying the character, we proceed to generate the final description. At this stage, the video, cast list, and subtitles with character names are provided to the model. Audio input is excluded to ensure compatibility with most existing open-source or closed-source large vision-language models. For open-source models, we use video description data from the MovieStory101 training set for supervised fine-tuning. During this stage, both the visual adapter and LLM are trained.
@@ 4 MovieStory101 Dataset
Figure 3: An example of the data in MovieStory101. Each consists of a 3-minute video, a detailed description, a corresponding cast list, and subtitles with character names.
@@
MovieNet [18] movie - 1100 42K 428.0 - -
ActivityNet Captions [22] ActivityNet activity description - 20k 180.0 49.2 -
YouCook2 [52] Youtube cook video activity description - 2k 316.2 67.8 -
ViTT [17] Youtube video activity tag - 8k 282.0 21.1 -
CMD-AD [15] movie audio description 1432 101k 2.8 9.1 -
Movie101 [49] movie audio description 101 14.1k 20.4 80.7 2.0
Movie101-v2 [50] movie audio description 203 46k 12.8 39.1 1.9
@@
MovieNet [18]
MovieQA [42]
CMD-AD [15]
Movie101 [49]
Movie101-v2 [50]
@@
We introduce MovieStory101, a long video description dataset based on movies. We choose movies because they contain rich and diverse information. A high-quality movie description should capture not only fine-grained details such as scenes, characters, and actions but also higher-level aspects like character relationships, plot development, and causal links between events (see Figure. 3).
The MovieStory101 task is partially related to audio description for movies [13, 14, 15], which narrates non-dialogue segments, and dense video captioning [22, 6], which generates temporally localized descriptions. However, MovieStory101 offers more comprehensive, plot-level description for long videos, integrating both visual and audio information. We refer to this as full description. In Table 1, we report the statistics of MovieStory101 and compare them with those of the other video description datasets, emphasizing its unique features: detailed descriptions, long video durations, and a large number of characters. Additionally, we provide a comparison of the annotations across different movie datasets in Table 2.
@@ D. Prepares for surgery
Action What does Jiang Feng do before going to bed at night? A. Drinks water
C. Takes a sleeping pill
B. Writes in a diary
D. Takes a shower
Character Who suggests Jiang Feng to rethink about his decision to undergo the memory erasing surgery? A. Zhang Daichen
C. Hui Lan
B. The boss
D. The person on the other side of the phone
@@ D. He feels relieved
Plot What must happen within 72 hours of the first memory reload for Jiang Feng? A. He needs another memory extraction surgery.
C. He needs to restore all his memories.
B. He must take a special pill.
D. He should make a reservation.
Plot Why does Hui Lan unplug the phone line? A. To prevent Jiang Feng from making a call
C. As per Jiang Feng’s suggestion
B. Because she is scared
D. To charge her phone
@@ 4.1 Movie Source
We build MovieStory101 based on Movie101 [49] and Movie101v2 [50]. Together, they include 203 movies, providing audio descriptions for non-dialogue scenes.
Since the movies in Movie101 and Movie101v2 contain audio descriptions integrated into the audio tracks, we collect the original editions to obtain the original audio tracks. After review, we select 187 movies for our dataset, excluding 16 movies because their original editions differed significantly from their barrier-free versions. Each selected movie is segmented into 3-minute clips. The final MovieStory101 dataset includes 5,210 / 140 / 632 clips in the training / development / testing split, totaling 300 hours. Each split contains video clips from different movies.
@@ 4.2 Annotation Process
The annotation process comprises three steps: movie descriptions, character photos, and audio-visual character identification. Figure 4 shows the statistics for descriptions and characters in MovieStory101.
Movie description. We annotate each movie clip with detailed description. To improve efficiency, annotators use audio descriptions and ASR data from Movie101 and Movie101v2 as references. Annotators are tasked with providing the descriptions of each event in the video, along with the corresponding start and end timestamps. The descriptions should include details about scenes, characters, actions, dialogues, and plot development. Figure 3 provides an example in MovieStory101 dataset. See annotation guidelines in Appendix A. In addition, annotators revise ASR data to generate subtitles.
Character photos. Annotators identify main characters and select a face for each in each movie clip, creating a cast list with names and corresponding faces (see Figure 3).
Audio-visual character identification. Annotators identify the speaker for each dialogue line in the clips. If the speaker is in the cast list, their character name is assigned; otherwise, a descriptive name is used. Figure 3 provides an example of audio-visual character identification.
@@ 4.3 Evaluation of Description
Evaluating long video descriptions is challenging due to the numerous details they contain, which increases evaluation costs. To address this, we develop an automatic evaluation method called StoryQA for the MovieStory101 test set. On average, each 3-minute video in StoryQA is accompanied by 38 multiple-choice questions focusing on visual actions, character relationships, and plot development. During evaluation, we employ GPT-4 [31] to answer these questions given the long description as context, and the QA accuracy serves as an automated metric for assessing the quality of the description. The prompt template is shown in Appendix F. Table 3 shows examples in StoryQA.

### v3

@@ Abstract
Existing large vision-language models (LVLMs) are largely limited to processing short, seconds-long videos and struggle with generating coherent descriptions for extended video spanning minutes or more. Long video description introduces new challenges, such as consistent character identification and plot-level descriptions incorporating both visual and audio information. To address these, we figure out audio-visual character identification, matching character names to each dialogue, as a key factor. We propose StoryTeller, a system for generating dense descriptions of long videos, incorporating both low-level visual concepts and high-level plot information. StoryTeller uses a multimodal large language model that integrates visual, audio, and text modalities to perform audio-visual character identification on minute-long video clips. The results are then fed into a LVLM to enhance consistency of video description. We validate our approach on movie description tasks and introduce MovieStory101, a dataset with dense descriptions for three-minute movie clips. To evaluate long video descriptions, we create StoryQA, a large set of multiple-choice questions for MovieStory101 test set. We assess descriptions by inputting them into GPT-4 to answer these questions, using accuracy as an automatic evaluation metric. Experiments show that StoryTeller outperforms all open and closed-source baselines on StoryQA, achieving 9.5% higher accuracy than the strongest baseline, Gemini-1.5-pro, and demonstrating a +15.56% advantage in human side-by-side evaluations. Additionally, incorporating audio-visual character identification from StoryTeller improves the performance of all video description models, with Gemini-1.5-pro and GPT-4o showing relative improvement of 5.5% and 13.0%, respectively, in accuracy on StoryQA.
Codehttps://github.com/hyc2026/StoryTeller
@@ 1 Introduction
Generating detailed video descriptions is a core challenge in video understanding. While large vision-language models (LVLMs) have made significant progress (Wang et al. 2024; Lin et al. 2024; Chen et al. 2024b; Wang, Yuan, and Zhang 2024), they remain limited to processing seconds-long videos and generating low-level visual concepts like objects, scenes and atomic actions. Real-world videos, such as movies, are much longer and require high-level information for effective description, such as plot development. Recent research works address long videos by segmenting them, generating descriptions for each segment, and combining them using large language models (Zhang et al. 2024; Han et al. 2023b, a). However, these methods struggle to preserve character identity and overlook audio cues, making coherent, detailed, plot-level descriptions difficult to achieve.
Thus, accurate character identification and tracking across both visual and audio modalities are essential for high-quality long video descriptions. While existing studies (Raajesh et al. 2024; Ji et al. 2024) have improved the quality of video descriptions by tracking character identification via visual cues, a critical gap remains in linking dialogue with on-screen characters to identify speakers. Bridging this gap is essential for enhancing long video descriptions.
@@
A major challenge in long video description is the lack of training and test data. To address this, we introduce MovieStory101, a new dataset focused on movies—long videos with high information density. The videos in MovieStory101 are from Movie101(Yue et al. 2023) and Movie101v2 (Yue et al. 2024), comprising 187 movies split into 5,982 three-minute clips. Each clip is annotated with detailed storyline descriptions, subtitles, and character identity labels for each dialogue line. The storyline annotations ensure a complete understanding of the movie’s plot. By training on this dataset, we aim to unleash the model’s ability to generate plot-level descriptions. In addition, evaluating long video descriptions is also challenging. Besides human evaluation, we propose an automated evaluation method for MovieStory101’s test set: StoryQA. StoryQA has 11 multiple-choice questions in average for each video, covering visual actions, character relationships, and plot details. We evaluate a long video description by inputting it into GPT-4 (OpenAI 2023) to answer these questions, using the accuracy as an evaluation metric.
Based on StoryQA and human side-by-side evaluation, we demonstrate that StoryTeller outperforms Gemini-1.5-pro (Reid et al. 2024), GPT-4o (OpenAI 2024), and several advanced open-source LVLMs (Lin et al. 2024; Li et al. 2024; Wang et al. 2024; Chen et al. 2023) in generating descriptions for long videos. Specifically, the 7B model-based StoryTeller achieves an accuracy in the StoryQA evaluation that is 9.5% higher than the best-performing baseline model, Gemini-1.5-pro. In human side-by-side evaluation, StoryTeller shows an advantage of +15.56% over Gemini-1.5-pro. Additionally, we verify the significant role of accurate audio-visual character identification on improving description. Incorporating audio-visual character identification results from StoryTeller as input consistently improves the performance of various LVLMs, with Gemini-1.5-pro and GPT-4o achieving relative accuracy gains of 5.5% and 13.0% on StoryQA, respectively.
@@ 2 Related Work
Long Video Description. Limited by the number of input frames, LVLMs are primarily effective for short video clips and typically generate descriptions focused on low-level visual concepts (Wang et al. 2024; Lin et al. 2024; Chen et al. 2024b; Wang, Yuan, and Zhang 2024; Lin et al. 2023; Li et al. 2023b). There remains a gap in describing longer videos, which may span several minutes or more, and in capturing higher-level content like character interactions and plot progression.Some research (Zhang et al. 2024; Han et al. 2023b, a) treats long videos as a series of short clips, incorporating previous descriptions when describing the current one, but this can lead to cumulative errors that affect consistency and coherence of the overall narrative (Han et al. 2023b). Another approach, Video ReCap (Islam et al. 2024), offers a recursive model for varying video lengths. Nevertheless, for longer videos, this method tends to produce summaries, limiting its ability to generate detailed descriptions.
@@
Audio-Visual Character Identification. Character identification is essential for high-quality video descriptions (Raajesh et al. 2024; Ji et al. 2024), particularly to maintain consistency across clips. While prior research has focused on visual-level identification (Nagrani and Zisserman 2018; Tapaswi, Bäuml, and Stiefelhagen 2012), leveraging audio cues to determine which character is speaking can straightforwardly improve descriptions. A closely related area of research is audio-visual diarization, which integrates both visual and auditory features to segment speakers in audio (Xu et al. 2022; He, Du, and Lee 2022). Nonetheless, current methods are still unable to identify the speakers’ identities.
Figure 1: Example from MovieStory101: Each sample includes a 3-minute video, a detailed description, cast list, and subtitles with character names. We compare the full description with audio description (from Movie101); green text highlights additional details in full description, such as character identification, emotions, and activities.
@@ 3 MovieStory101 Dataset
MovieNet (Huang et al. 2020b) movie - 1100 42K 428.0 - -
ActivityNet Captions (Krishna et al. 2017) ActivityNet activity description - 20k 180.0 49.2 -
YouCook2 (Zhou, Xu, and Corso 2018) Youtube cook video activity description - 2k 316.2 67.8 -
ViTT (Huang et al. 2020a) Youtube video activity tag - 8k 282.0 21.1 -
CMD-AD (Han et al. 2024) movie audio description 1432 101k 2.8 9.1 -
Movie101 (Yue et al. 2023) movie audio description 101 14.1k 20.4 80.7 2.0
Movie101-v2 (Yue et al. 2024) movie audio description 203 46k 12.8 39.1 1.9
@@
MovieNet (Huang et al. 2020b)
MovieQA (Tapaswi et al. 2016)
CMD-AD (Han et al. 2024)
Movie101 (Yue et al. 2023)
Movie101-v2 (Yue et al. 2024)
@@
MovieStory101 is a long video description dataset based on movies from Movie101 (Yue et al. 2023) and Movie101v2 (Yue et al. 2024). We choose movies because they contain rich and diverse information. A high-quality movie description should capture not only fine-grained details such as scenes, characters, and actions but also higher-level aspects like character relationships, plot development, and causal links between events (see Figure 1).
The MovieStory101 task is related to audio description for movies (Han et al. 2023b, a, 2024), which narrates non-dialogue segments, and dense video captioning (Krishna et al. 2017; Chen et al. 2020), which generates temporally localized descriptions. However, MovieStory101 goes further by offering comprehensive, plot-level description, integrating both visual and audio information. We refer to this as full description. Figure 1 illustrates the difference between a full description and the original audio description from Movie101. In Table 1, we report the statistics of MovieStory101 and compare them with those of the other video description datasets, emphasizing its unique features: detailed descriptions, long video durations, and a large number of characters. Additionally, we provide a comparison of the annotations across different movie datasets in Table 2.
@@ D. He feels relieved
Plot What must happen within 72 hours of the first memory reload for Jiang Feng? A. Another memory extraction surgery
C. Restore all his memories
B. Take a special pill
D. Make a reservation
@@ 3.1 Movie Source
We construct MovieStory101 based on Movie101 and Movie101v2, which together comprise 203 movies. Since both sources contain audio descriptions integrated into the audio tracks, we collect the original editions to obtain the original audio tracks. After review, we select 187 movies for our dataset, excluding 16 movies because their original editions differed significantly from their barrier-free versions. Each selected movie is segmented into 3-minute clips. The final MovieStory101 dataset includes 5,210 / 140 / 632 clips in the training / development / testing split, totaling 300 hours. Each split contains video clips from different movies.
@@ 3.2 Annotation Process
The annotation process includes movie descriptions, character photos, and audio-visual character identification.
Movie description. We annotate each movie clip with detailed description. To improve efficiency, annotators use audio descriptions and ASR data from Movie101 and Movie101v2 as references. Annotators are tasked with providing the descriptions of each event in the video, along with the corresponding start and end timestamps. The descriptions should include details about scenes, characters, actions, dialogues, and plot development. In addition, annotators revise ASR data to generate subtitles.
Character photos. Annotators identify main characters and select a face for each in each movie clip, creating a cast list with names and corresponding faces.
Audio-visual character identification. Annotators identify the speaker for each dialogue line in the clips. If the speaker is in the cast list, their character name is assigned; otherwise, a descriptive name is used.
See annotation guidelines and annotator information in Appendix A. Figure 1 provides an example in MovieStory101 dataset. Figure 2 shows the statistics for descriptions and characters in MovieStory101.
@@ 3.3 Evaluation of Description
Evaluating long video descriptions is challenging due to the numerous details they contain, which increases evaluation costs. To address this, we develop an automatic evaluation method called StoryQA for the MovieStory101 test set. On average, each 3-minute video in StoryQA is accompanied by 38 multiple-choice questions focusing on visual actions, character relationships, and plot development. During evaluation, we employ GPT-4 (OpenAI 2023) to answer these questions given the long description as context, and the QA accuracy serves as an automated metric for assessing the quality of the description. The prompt template is shown in Appendix F. Table 3 shows examples in StoryQA.
Figure 4: Pipeline of StoryTeller, consisting of three modules: Stage I - Video Segmentation, where videos are divided into short clips that are relatively independent and internally complete; Stage II - Audio-Visual Character Identification, where characters are identified for each dialogue line using both audio and visual cues; and Stage III - Description Generation, where detailed descriptions are generated for each short clip, ultimately producing a consistent narrative for the entire long video.
@@
4 StoryTeller Long Video Description System
4.1 Overview
4.2 Audio-Visual Character Identification
The audio-visual character identification module determines speakers identity in video clips by integrating audio, visual and textual information. In particular, given a video clip (frames, audio and subtitles) and a cast list (character photos and names), the module assigns each dialogue line to its corresponding character. If the speaker is in the cast list, their name is provided; otherwise, a descriptive identity (e.g., ‘a police officer’ or ‘a little girl’) is used. This task presents several challenges: 1) relying solely on visual information is often insufficient, especially when multiple characters are on screen or the speaker is not visible; 2) audio cues in seconds-long clips are typically isolated, making them less effective for character identification. However, in longer segments, audio can be leveraged to extract speaker features across different clips, such as identifying if the same person speaks in multiple segments. Models that are limited to short segments often struggle to incorporate such global features effectively.
To address the aforementioned challenges, we design a novel audio-visual character identification mechanism incorporating a MLLM equipped with both audio and video modalities. We also develop a global decoding algorithm to leverage global audio information for more accurate character identification. Specifically, as illustrated in Figure 4, we begin by performing audio diarization on the entire video, segmenting the dialogue, and assigning IDs to different speakers (e.g., C1, C2, etc.), as detailed in Appendix D. The subtitles with global IDs are then fed into the MLLM to identify the corresponding character names for each global ID. The global decoding mechanism processes short video clips in parallel, ensuring that global IDs in different clips are consistently mapped to the same name, thereby improving identification accuracy. Details are provided below.
Model Architecture and Training. We integrate a MLLM capable of processing both visual and auditory data, as depicted in Figure 5. For visual inputs, each image or video frame is encoded by CLIP-ViT (Radford et al. 2021), followed by mapping into the token embedding space of the downstream LLM through a multi-layer perceptron (MLP). These tokens are then fed into the LLM. The audio inputs are processed by dual auditory encoders (Tang et al. 2023), a speech encoder from Whisper-Large-v2 (Radford et al. 2023) and a non-speech BEATs audio encoder (Chen et al. 2022). The outputs from the two audio encoders are concatenated at the frame level. Subsequently, a window-level Q-former (Li et al. 2023a) align the combined audio representation to the LLM’s input space.
We initialize the visual module and language model from the Tarsier-7b (Wang, Yuan, and Zhang 2024), a large vision-language model demonstrating SOTA performance on various video understanding benchmarks. Our training protocol consists of three phases: (1) pre-training the audio module, (2) fine-tuning on tasks including audio diarization, character identification, recognition and tracking, and (3) fine-tuning for audio-visual character identification using MovieStory101 training set. Details on the dataset used in each phase and the complete training methodology are provided in Appendix E.
Global Decoding. The MLLM assigns names to each global character ID. First, these names come from a set $\mathcal{N}$, including all names from the cast list as well as a special item “Others”, representing characters not in the cast list. When the model outputs “Others”, it will further provide a descriptive name. Second, the model outputs a JSON dictionary, where each key is a global character ID within the input short clip, and each corresponding value is the assigned name. Keys are listed in ascending order of global character IDs.
$p_{\theta}(y|v_{k},x)=\prod_{t}p_{\theta}(y_{t}|v_{k},x,y_{<t}),$ (1)
Model StoryQA Accuracy $\uparrow$ CIDEr $\uparrow$ R@1/5 $\uparrow$ CRITIC $\uparrow$
Character Action Plot Total
Gemini-1.5-pro 0.578 0.501 0.534 0.544 0.035 0.231 0.497
GPT-4o 0.517 0.479 0.528 0.507 0.000 0.318 0.370
VILA1.5-8B 0.561 0.459 0.540 0.524 0.120 0.339 0.608
LLaVA-OneVision-7B 0.557 0.454 0.540 0.520 0.199 0.339 0.615
Qwen2-VL-7B 0.549 0.468 0.549 0.523 0.084 0.340 0.594
InternVL2-8B 0.535 0.448 0.506 0.501 0.168 0.336 0.573
StoryTeller 0.676 0.583 0.644 0.639 0.237 0.342 0.630
Table 4: Results of video descriptions generated by StoryTeller and baselines, evaluated using StoryQA and traditional metrics.
Then, the global probability of $x$ corresponding to name $y$ is defined as
$p^{\rm global}_{\theta}(y|x)=\prod_{k=1}^{m}p_{\theta}(y|v_{k},x).$ (2)
For example, as shown in Figure 5, both videos $v_{1}$ and $v_{2}$ contain lines from character $C1$. In $v_{2}$, although Xu Tailang speaks, the camera focuses entirely on silent Director Jin, leading the model assigns a higher probability to Director Jin being the speaker. However, the model still assigns over 30% confidence to Xu Tailang. By integrating information from the entire video globally, $C1$ is ultimately mapped to Xu Tailang, correcting the erroneous prediction in $v_{2}$.
4.3 Description Generation
After identifying the character, we proceed to generate the final description. At this stage, the video, cast list, and subtitles with character names are provided to the model. Audio input is excluded to ensure compatibility with most existing open-source or closed-source large vision-language models. For open-source models, we use video description data from the MovieStory101 training set for supervised fine-tuning. During this stage, both visual adapter and LLM are trained.
Figure 6: Comparison of StoryTeller and Gemini-1.5-pro on a video example featuring multiple characters (same as Figure 1).